OpenStudy (cooltowl):

f(x)=(x+4)^2-9 determine were the f(x) intersects the x-axis

OpenStudy (anonymous):

intercepting with the x axis means f(x)=0

OpenStudy (anonymous):

\[0=(x+4)^2 -9\]

OpenStudy (anonymous):

\[0=x^2+8x+7\]

OpenStudy (anonymous):

\[(x+1) (x+7) = 0\]

OpenStudy (anonymous):

\[x= -1 x= -7\]

OpenStudy (anonymous):

\[y'=2x+8\]

OpenStudy (anonymous):

\[y''=2\]

OpenStudy (anonymous):

meaning minimum turning point so if you were to plot this it should have a minimum turning point and pass through those points

OpenStudy (anonymous):